OpenStudy (anonymous):

Macy is feeding the pigeons at her local park. The number of pigeons around her doubles every 15 minutes. The following function represents the number of pigeons around her after x minutes. P(x)=5(2^x/15) Which statement is true? The expression, (0.133)^x, reveals the approximate rate of increase of the pigeons per hour. The expression, (1.047)^x, reveals the approximate rate of increase of the pigeons per minute. The expression, (0.133)^x, reveals the approximate rate of increase of the pigeons per minute. The expression, (1.047)^x, reveals the approximate rate of increase of th

OpenStudy (anonymous):

Use the properties of exponents to rewrite the function. P(x)=5(2^x15) =5(2^1/15)^x =5(1.047)^x In the expression, 5 represents the initial number of pigeons, and (1.047)^x reveals the rate of increase. Therefore, the expression, (1.047)^x, reveals the approximate rate of increase of the pigeons per minute.

OpenStudy (anonymous):

so its B